Find the potential at a distance of 0.07 m from a -6 nC charge

1
Expert's answer
2022-03-08T17:00:08-0500

Potential

"V=\frac{kq}{r}"


"V=\frac{-9\times10^9\times6\times10^{-9}}{0.07}=-771.42Volt"